# Apple Vision Pro: Comments in Healthcare

### v1

@@ Abstract
This paper objectively analyzes the emerging discourse surrounding Apple Vision Pro’s application in healthcare and medical education. Released in June 2023, Apple Vision Pro represents a significant advancement in spatial computing, combining augmented and virtual reality to create new possibilities in digital interaction. We aim to compile and present recent articles. We used PubMed, IEEE Xplore, Google Scholar, and JSTOR. Non-academic publications were excluded. The results were six commentaries, one a pre-print. All were majorly optimistic, with one mentioning VR/AR sickness. For future research directions, we stress the need for continued exploration of Apple Vision Pro’s capabilities and limitations and expect expert opinions to englobe this discussion.
Index Terms:
Apple Vision Pro, Apple visionOS
I Introduction
I-A Context Setting
@@ I-B Purpose of the Discussion
II Methodology
Using academic databases, including PubMed, IEEE Xplore, Google Scholar, and JSTOR, we used the keywords: ”Apple Vision Pro” and ”Apple visionOS”.
II-A Criteria for Inclusion and Exclusion of Papers
@@
  • Papers published post-June 2023 for up-to-date relevance.
  • Works presenting empirical data, theoretical analysis, or relevant commentary on the technology.
@@
  • Papers not focusing on Apple Vision Pro or Apple visionOS.
@@
II-B Results
We identified six papers that provided diverse perspectives on Apple Vision Pro. This selection comprised comment articles, including one pre-print from arXiv authored by a team with relevant academic expertise. Although the pre-print was not published, its inclusion was justified by the authors’ knowledge and insights into Apple Vision Pro’s applications.
Our goal was to cover a range of viewpoints on the technology, ensuring a balanced representation of current state contributions from recognized sources.
III Expected Applications of Apple Vision Pro in Healthcare and Medical Education
The key points from each commentary are summarized in Table I, offering a concise overview of the perspectives on Apple Vision Pro’s applications in these critical fields.
III-A Healthcare: Precision Medicine and Clinical Procedures
@@
III-B Medical Education and Training
@@
III-C Ophthalmology and Vision Science
@@
III-D Enhancing Accessibility in Healthcare
@@ III-E Psycological Research
III-F Advancement in Medical Education
@@
TABLE I: Summary of Results
@@
Egger et al. (2023)
Healthcare
Clinical procedures, patient care
Enhanced surgical navigation, improved patient interaction
Waisberg et al. (2023, Eye)
Ophthalmology
Vision screening, diagnostics
Precise diagnostics, improved patient education
Waisberg et al. (2023, Irish Journal)
Medical Education
Training for medical professionals
Interactive learning, improved training methodologies
Masalkhi et al. (2023)
Healthcare Accessibility
Enhanced visual experiences for the visually impaired
Improved accessibility in healthcare settings
Barrett et al. (2023)
Psychological Research
VR-based emotional and behavioral studies
Enhanced research capabilities, realistic 3D rendering
Waisberg et al. (2023, Canadian Journal)
Medical Education
Advanced medical training
Immersive educational experiences enhanced diagnostic skills
IV Discussion
IV-A Challenges and Future Directions
While the papers reviewed paint an optimistic picture of Apple Vision Pro’s applications [2] [3] [4] [5] [6] [7], they also acknowledge various challenges and future directions. These include integrating the technology seamlessly into existing systems (especially in healthcare), addressing privacy and data security concerns, and ensuring the device’s accessibility and user adaptability (e.g., in a clinical setting [5]). The current high price point could limit accessibility, especially in resource-constrained healthcare settings or educational institutions, although in some settings, this might not be an issue[2].
IV-B VR/AR Side Effects
@@
V Final Thoughts

### v4

@@ Abstract
This paper objectively analyzes the emerging discourse surrounding the application of Apple Vision Pro in healthcare and medical education. Released in June 2023, Apple Vision Pro represents a significant advancement in spatial computing by combining augmented and virtual reality to create new possibilities for digital interaction. We aim to compile and present recent scholarly articles, utilizing PubMed, IEEE Xplore, Google Scholar, and JSTOR as our primary sources. Non-academic publications were excluded. The research yielded six commentaries, including one pre-print and one case report, primarily expressing optimism about the technology, albeit with some concerns regarding VR/AR sickness. We highlight the need for ongoing exploration of Apple Vision Pro’s capabilities and limitations and anticipate that expert opinions will further enrich this discourse.
keywords:
Apple Vision Pro, Apple Vision
1 Introduction
1.1 Context Setting
@@ 1.2 Purpose of the Discussion
2 Methodology
We conducted a comprehensive literature review using academic databases such as PubMed, IEEE Xplore, Google Scholar, and JSTOR. The keywords ”Apple Vision Pro” and ”Apple visionOS” were employed to identify relevant publications.
2.1 Criteria for Inclusion and Exclusion of Papers
@@
  • Papers published after June 2023 to ensure up-to-date relevance.
  • Works presenting empirical data, theoretical analysis, or significant commentary on the technology.
@@
  • Papers not directly focusing on Apple Vision Pro or Apple visionOS.
@@
2.2 Results
Our search yielded six papers that provided diverse perspectives on Apple Vision Pro, including comment articles, one pre-print from arXiv, and one case report. The inclusion of the pre-print was justified by its authors’ relevant academic expertise and insights into the applications of Apple Vision Pro, enriching the discourse on its potential impact in healthcare and medical education.
3 Expected Applications of Apple Vision Pro in Healthcare and Medical Education
The key points from each commentary are summarized in Table 1, offering a concise overview of the perspectives on Apple Vision Pro’s applications in these critical fields.
3.1 Healthcare: Precision Medicine and Clinical Procedures
@@
3.2 Medical Education and Training
@@
3.3 Ophthalmology and Vision Science
@@
3.4 Enhancing Accessibility in Healthcare
@@ 3.5 Psycological Research
3.6 Advancement in Medical Education
@@
3.7 Neurosurgical Planning Tool
Olexa et al. (2024): The Apple Vision Pro as a Neurosurgical Planning Tool: A Case Report.[8] This case report demonstrates the practical application of the Apple Vision Pro in neurosurgery, specifically in the planning and execution of a ventriculoperitoneal shunt placement for a patient with hydrocephalus. Using the Apple Vision Pro headset, surgeons were able to visualize and interact with a 3D model of the patient’s anatomy, significantly aiding in the surgical planning process. The technology allowed for an enhanced understanding of the anatomical structures involved, particularly the ventricles’ size and location. Following the procedure, clinicians provided feedback via a survey, reporting that the 3D models felt realistic and the integration of the user’s real-world view through the headset felt natural, without causing eye strain or fatigue. Importantly, the majority expressed a willingness to continue using this technology for future cases. This case report highlights the potential of the Apple Vision Pro to revolutionize neurosurgical planning and anatomical visualization, offering an example of its real-world application in a clinical setting and underscoring the promise it holds for improving surgical precision and patient outcomes as the device becomes more integrated into healthcare practices.
Table 1: Summary of Results
@@ 3.7 Neurosurgical Planning Tool
Egger et al. (2023) Healthcare Clinical procedures, patient care Enhanced surgical navigation, improved patient interaction
Waisberg et al. (2023, Eye) Ophthalmology Vision screening, diagnostics Precise diagnostics, improved patient education
Waisberg et al. (2023, Irish Journal) Medical Education Training for medical professionals Interactive learning, improved training methodologies
Masalkhi et al. (2023) Healthcare Accessibility Enhanced visual experiences for the visually impaired Improved accessibility in healthcare settings
Barrett et al. (2023) Psychological Research VR-based emotional and behavioral studies Enhanced research capabilities, realistic 3D rendering
Waisberg et al. (2023, Canadian Journal) Medical Education Advanced medical training Immersive educational experiences, enhanced diagnostic skills
Olexa et al. (2024) Neurosurgery Surgical planning for ventriculoperitoneal shunt placement Enhanced precision in surgical planning, improved anatomical visualization
4 Discussion
4.1 Challenges and Future Directions
The collection of papers reviewed, including commentaries and a clinical case report [2] [3] [4] [5] [6] [7] [8], presents a predominantly optimistic view of the potential applications of Apple Vision Pro in healthcare and medical education. However, they also shed light on various challenges that must be addressed.
Integrating Apple Vision Pro into existing healthcare systems poses a significant challenge, as illustrated by the case report [8], which underscores the device’s potential in neurosurgical planning and execution. While the device demonstrated considerable promise in enhancing surgical precision and patient outcomes, its adoption is hindered by issues such as the need for seamless integration into hospital workflows, privacy and data security concerns, and ensuring the device’s accessibility and adaptability for users in clinical settings [5].
The case report by Olexa et al. (2024) [8] not only exemplifies the practical application of Apple Vision Pro in a complex neurosurgical procedure but also highlights the possibilities of use cases for continued research and development.
Moreover, the high cost of Apple Vision Pro could limit its accessibility, particularly in resource-constrained settings [2]. This financial barrier might restrict the technology’s widespread adoption, emphasizing the need for cost reduction strategies without compromising the device’s quality and functionality.
In conclusion, while the Apple Vision Pro stands on the cusp of revolutionizing healthcare and medical education through its advanced spatial computing capabilities, realizing its full potential requires a concerted effort to address these challenges. The insights provided by Olexa et al. (2024) offer valuable perspectives for future endeavors aimed at integrating this innovative technology into mainstream medical practice.
4.2 VR/AR Side Effects
@@
4.3 Apple Vision Pro Health Advisory
In a report dated January 19, 2024, Apple Inc. issued a health advisory for users of Apple Vision Pro, particularly those with certain medical conditions such as migraines or chronic headaches, dizziness or vertigo, eye or vision conditions, such as binocular vision conditions, psychological conditions, inner ear conditions, history of dry eyes, itchiness, or swelling of the eyelids, skin allergies or sensitivities, seizures, balance or gait conditions[9]. The company recommends these individuals consult with a healthcare provider before using the device. Also, updated acknolegement to when to stop using Apple Vision Pro: if you experience symptoms related to a medical condition; if your medical provider confirmed that it’s safe for you to use Apple Vision Pro, but you experience severe or persistent physical discomfort, motion sickness, visual discomfort, skin irritation, or psychological symptoms; if you notice swelling, itchiness, skin irritation, or other skin reactions while using your device or after using the device[9]. Always, if symptoms persist, consult with your medical provider[9].
This advisory highlights the importance of considering individual health conditions when using advanced technological devices in medical settings.
5 Final Thoughts
@@
Acknowledgment
The authors declare that there is no conflict of interest regarding the publication of this paper. This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.